# HateCOT: An Explanation-Enhanced Dataset for Generalizable Offensive Speech Detection via Large Language Models

### v2

@@ Abstract
Warning: This paper contains examples of very offensive content. The ubiquitousness of social media has led to the need for reliable and efficient detection of offensive content to limit harmful effects, leading to the proliferation of datasets and models related to detecting offensive content. While sophisticated models have attained strong performance on individual datasets, they often do not generalize due to differences between how “offensive content” is conceptualized, and how these datasets are labeled. In this paper, we introduce HateCOT, a dataset of 52,000 samples drawn from diverse existing sources with explanations generated by GPT-3.5-Turbo and human-curated. We show that pre-training models for the detection of offensive content on HateCOT significantly boots open-sourced Language Models on three benchmark datasets in both zero and few-shot settings, despite differences in domain and task. We further find that HateCOT enables effective K-shot fine-tuning in low-resource settings.111 Our repository is available at [REDACTED]
@@ 1 Introduction
As social media has become an indispensable component of modern discourse, this channel of communication has also amplified the propagation of offensive content. Speech aimed to promote hateful sentiments thrive on such platforms, leading to real and significant harm on their audience Giachanou and Rosso (2020); Saha et al. (2019). However, “offensive content” is an essentially contested construct, and exactly what is and is not allowed varies by platform. In research, different approaches analyze semantically similar but still distinct concepts: Cyber-bullying, Toxicity, Sexist, Racist, Hate etc. Poletto et al. (2021); Fortuna et al. (2021); Nghiem et al. (2024), making visible this contestedness.
Compounding the challenge, reliable detection of offensive content requires significant data resources. Sophisticated models tend to be data-hungry, and the process of curating a dataset tailored to specific use cases can be costly, time-consuming, and emotionally challenging for annotators Founta et al. (2018); Toraman et al. (2022). The typical pipeline consists of collecting samples based on topic-relevant key words, then recruiting either crowdworkers or experts to annotate data before developing classification models Paullada et al. (2021). Each step incurs investment and may inject subtle idiosyncrasies proportionate to the size of the downstream dataset, further limiting potential transferable usefulness to related tasks Fortuna et al. (2021). Furthermore, the size of a dataset does not necessarily guarantee cross-main transferrability Poletto et al. (2021); Fortuna et al. (2021).
In practical settings, users may desire transparency from social media platforms. Therefore, the ability to provide human understandable justification based on platform-specific policy becomes an attractive feature for content moderation. Nevertheless, current techniques often still fail to offer intuitive explanatory signals, especially to end users Yadav et al. (2023); Babaeianjelodar et al. (2022); Ibrahim et al. (2022).
@@
    We curate and release HateCOT (Hate-related Chains-of-Thought) a dataset of $52k$ samples consisting of input text, a hate speech label, and an explanation of that label. We collect this data by merging eight datasets with explanations created by using GPT-3.5-Turbo to augment human annotations.
@@
    We demonstrate the efficacy of using HateCOT as a pre-training corpus to enable finetuning on a target domain and task. We show that open-sourced LLMs can effectively leverage definitions to adapt to new tasks using zero-shot and few-shot settings via finetuning and in-context learning across 3 datasets.
@@
    We assess the quality of explanations generated by our finetuned models with respect to the criteria described in their corresponding dataset’s definitions. These insights showcase LLM-generated explanations as a means to enhance transparency in content moderation.
@@ 2.1 Offensive Speech Detection
Cross-domain generalization, however, still remains a relevant challenge in the area. Fortuna et al. (2021) found empirically that cross-dataset transference is highly dependent on semantic similarity between their label spaces. Recent works have pre-trained Transformer-based models, such as HateBERT and fBERT, on specialized corpora to enhance generalization to various levels of success Caselli et al. (2021); Sarkar et al. (2021).
@@ 2.2 LLMs in Offensive Speech Classification
Zampieri et al. (2023) assessed a range of open-sourced LLMs on zero-shot prompting on the OffensEval task and found their performance trailing by a wide margin behind trained existing fine-tuned BERT-based systems. Chiu et al. (2021) and Han and Tang (2022) used the proprietary GPT-3 on a set of different datasets and noted that informative contexts and examples could boost the model’s performance to competitive levels on a different set of data. Similarly, Roy et al. (2023) found that adding explanation to pipeline could result in 10 to 20% boost in performance of LLMs over baselines. (Yang et al., 2023)’s study found that training LLMs with step-by-step reasoning grounded by annotations could improve predictive power.
Pre-trained language models have exhibited remarkable ability in text generation Celikyilmaz et al. (2020). Recent large-size LLMs such as GPT-3 and later models are capable of astoundingly fluent, convincing and knowledge-infused outputs Zhang et al. (2023). LLMs with hundreds or more parameters even exhibit reasoning capabilities Wei et al. (2022, 2021), leading to a flurry of research on prompting techniques to harness their prowess, such as Chain-of-thought (COT), Tree-of-thought etc. (Yao et al., 2023; Diao et al., 2023). An interesting line of research leverages LLMs to efficiently generate high volumes of synthetic data for tasks with training resource is scarce Puri et al. (2020); Bao et al. (2023); Whitehouse et al. (2023). We build upon these works to construct a dataset that can induce smaller LLMs to efficiently adapt to new categories of offensive content by leveraging their provided definitions.
@@ 3.1 Data Selection
    Target / Rationale: the dataset should elicit the a) targets and/or b) rationales, which are fragments of free texts by human annotators to convey some understanding of the corresponding post. Figure 4 shows several examples of rationales, demonstrating their unsuitability to be used as explanations in their native format.
These criteria significantly reduce the number of eligible candidates since many do not contain information on target and the associated definitions. Table 1(a) lists the 8 selected datasets.
@@
Using similar criteria, we also select 3 datasets with different label spaces and definitions to test downstream models’ generalizability (details in Table 1(b)).
HateCheck was created by Röttger et al. (2021) with the explicit goal of evaluating hate speech detection models. The authors recruited and trained 10 annotators on label the cases on a binary schema: Hateful or Non-hateful, and reported an almost perfect final inter-annotator agreement of 0.93.
HateXplain was constructed by Mathew et al. (2021). The authors asked annotators to highlight the span of tokens, called rationales, that contribute to their selection of the labels: Hate, Offensive, Normal. This dataset is primarily collected from Twitter and the Gab platform.
Latent_Hate was created by ElSherief et al. (2021a) on the premise that offensive speech classifiers tend to bias towards covert negative sentiment. With 22, 584 samples, this dataset is collected from Twitter( we discard the augmentation from the Social_Bias dataset avoid contaminating the training corpus (Sap et al., 2019)) . This set contains 2 coarse-grained labels Not Hate, Explicit Hate, Implicit Hate, while a subset contains 6 fine-grained categories, which we refer to as Implicit_Hate.
@@
Yang et al. (2023) showed that using GPT-3 to augment human-written rationales creates linguistically coherent while maintains faithfulness to the original content. We use the template shown in Figure 1 to generate the explanation for each input guided by the available annotations on label, target, and rationale from the chosen corpora. For datasets that contain multiple annotations per sample, we select the ultimate label via majority voting and concatenate annotations on the targets and/or rationales into a single string delimited by "|".
@@
We choose GPT-3.5-Turbo accessed via the OpenAI’s API to generate explanations due to the balance between the model’s affordability and its capability to follow instructions and generate coherent outputs (Ye et al., 2023a; Koubaa, 2023). For each of the 8 training datasets, we first randomly select and qualitatively analyze 20 samples to ensure the generated explanations are a) stylistically coherent, b) consistent with the provided labels, and c) congruent with the criteria denoted by the definitions. If deemed unsatisfactory, we iteratively adjust the input prompt until the quality threshold is achieved. Appendix A.3 reveals more specific details on the final prompts for different scenarios.
@@ 3.2 Optimization of Synthesized Corpus
Extending previous works Magister et al. (2022); Ho et al. (2022), we are interested in optimizing 2 parameters central to the construction of our corpus: the distribution of neutral vs. non-neutral classes in the data and the number of explanations per sample. The former has been noted to influence predictive powers Rathpisey and Adji (2019); Casula and Tonelli (2020). The latter, also referred to as degree of reasoning diversity, could improve knowledge distillation Ho et al. (2022).
3.2.1 Setup
Below are the experiments we perform on a sample of collected data to optimize these 2 parameters.
@@ 3.2.1 Setup
Figure 5 illustrates the Alpaca-styled template to format each post with its corresponding label, generated explanation, definitions along with the instruction into blocks of an input prompt (Taori et al., 2023). Using the described corpus, we supervised finetune the open-sourced model Llama 2 Chat-HF of 7 billion parameters (hereby referred to as Llama 7B) (Touvron et al., 2023) via LoRA techniques (Technical details specified in Appendix A.2) (Hu et al., 2021). This model is selected as the primary representative for this stage due to its manageable size and strong classification performance Touvron et al. (2023). Then, we perform zero-shot classification using the same template to prompt the fine-tuned models to generate the explanation and label for posts drawn from the 3 test datasets. For this part, posts are drawn using K-shot sampling (equal number of samples for each class) on the Validation set of the test data (K val in Table 2(b)). We omit Implicit_Hate at this stage due to this set’s markedly different 6-label space.
@@
For the first experiment on the ratio of neutral vs. non-neutral categories, the training data is split into subsets with their respective distribution between the neutral (NE) class and non-neutral class(es) (NN) described by the following formula: $NN=R*NE$, where $R\in\{1,2,3,4\}$ is the ratio coefficient. We set the number of explanations per sample to 2 , the smallest value that still enables the benefit of reasoning diversity. For the second experiment, we construct the subsets by varying the degree of reasoning diversity $D\in\{1,2,3,4\}$, of each post.
@@
We extract the explanation and predicted labels after their respective tag in the generated output. If the models generate multiple items from the dataset’s label space, we select the first admissible label. If no acceptable output is obtained, we randomly select an item in the label space.
@@ 3.2.2 Insights and Augmentation
We report the model’s macro F1-scores on the validation set of each configuration in Table 2. A balanced distribution between neutral and non-neutral classes in the training corpus is beneficial, as reflected by the substantially high mean F1-score of 0.643 when R=1. On the other hand, having 3 explanations per sample (D=3) achieves the best overall performance across 3 test sets, consistent with Ho et al. (2022)’s findings on the benefit of multiple reasoning paths. Nevertheless, performance markedly degrades when D=4. Based on manual analysis, we hypothesize that the quality of generated outputs degrades as the degree of diversity increases, consequently affecting the performance of models trained on this data. From these results, we augment the training corpus by selecting an additional approximately 1,800 samples from each of the 8 datasets while preserving the 1:1 balanced ratio of neutral to non-neutral classes. Then we collect 3 explanations per sample using the described mechanism, resulting in a final corpus of 52,137 samples (Table 1(a)), hereby referred to as HateCOT.
@@ 4 Experiments on Test Sets
We perform experiments to answer 3 questions. First, does HateCOT improve zero-shot classification of open-sourced LLMs on unseen datasets? Second, how much data is necessary to enable competitive performance via in-domain fine-tuning after pretraining on HateCOT? Finally, is in-context learning a viable alternative to finetuning?
@@
    COT-T5-XL This model is a variant of the Flan-T5-XL (3 billion parameters) model that is further finetuned on the CoT dataset, a collection of 1.8 million samples augmented with chain-of-thought-style explanations (Kim et al., 2023).
@@ 4.1 Zero-shot Classification
This setting prompts models to perform the classification task with no in-context examples via 2 modes: No Explanation, which prompts the model to directly predict the label for the given input, and With Explanation, which requires a justification before the predicted label. We finetune the base models using only HateCOT and evaluate their performance on the 4 test sets as in Section 3.2.2 (technical details in Appendix A.2).
With results presented in Figure 2(a), we observe that the smaller Flan-T5-L and OPT-IML are unable to generate explanations when prompted. In contrast, their scaled-up counterparts could follow instructions at all settings. Asking Base (off-the-shelve) models to provide explanations before label provides observable boost to Llama models in HateCheck and HateXPlain datasets, but actually hampers performance in Latent_Hate and its derivative Implicit_Hate, datasets notably challenging due to its covert nature ElSherief et al. (2021b).
@@
Finetuning on HateCOT unanimously enables all models to generate explanations. While smaller models receive no observable boost in their performance, larger models (COT-T5-XL, Llama 7B, Llama 13B) are considerably enhanced compared to their base counterparts. For COT-T5-XL, with the exception of HateXplain, HateCOT-trained version attain F1 scores increment of 7.6% (HateXplain and 9.5% (Latent_Hate) over the base counterpart without explanations. Llama 7B observes 23.9%, 25.6%, 10% increment for HateCheck, HateXplain, Latent_Hate, respectively. These statistics are 27.9%, 118.5%, and 10.2% for Llama 13B. For Implicit_Hate, all models struggle to attain satisfactory performance. Llama models’ performance improve with HateCOT, but still not much better than chance.
The reduction in performance of HateCOT-pretrained models compared to their Base counterparts without explanations (Flan-T5-L and OPT-IML on almost all test sets, COT-T5-XL on HateXPlain and Implicit_Hate) is in line with literature as COT-style prompting tends to favor larger models Wei et al. (2022); Wang et al. (2024); Suzgun et al. (2023). These models’ impacted performance when asked for explanation is unsurprising considering their significantly smaller size and original pre-training data, especially in zero-shot settings. Iyer et al. (2022); Chung et al. (2022).
(a) Macro F-1 scores of LLMs in zero-shot setting using 3 configurations. Base refers to out-of-the-box models, +HateCOT denotes their finetuned counterpart on our dataset. FT5L: Flan-T5-L, CT5XL: COT-T5-XL. Results for Base Flan-T5-L and OPTIML models for With Explanation settings omitted to reflect their inability to generate explanation according to the prompt.
(b) Macro F-1 scores for models in zero-shot setting with explanation after K-shot in-domain finetuning at various values of K. Dashed line represents finetuned base models, solid line represents models pre-trained on HateCOT, then in-domain finetuned. For each dataset, the horizontal dashed line represents the base version of the best performing model at K=256 which is finetuned using the entire training data without any rationale for comparison, denoted as Best model (Full data, No. exp).
(c) Min, max and mean of macro F-1 scores for Llama 7B and Llama 13B over 5 seed using in-context learning. Pretrained denotes models finetuned on HateCOT only. Finetuned denotes ICL performed on models that have been both HateCOT-pretrained then K=256 shot in-domain finetuned.

### v4

# Instructions for *ACL Proceedings

@@ Abstract
Warning: This paper contains examples of very offensive material. The widespread use of social media necessitates reliable and efficient detection of offensive content to mitigate harmful effects. Although sophisticated models perform well on individual datasets, they often fail to generalize due to varying definitions and labeling of "offensive content." In this paper, we introduce HateCOT, an English dataset with over 52,000 samples from diverse sources, featuring explanations generated by GPT-3.5-Turbo and curated by humans. We demonstrate that pretraining on HateCOT significantly enhances the performance of open-source Large Language Models on three benchmark datasets for offensive content detection in both zero-shot and few-shot settings, despite differences in domain and task. Additionally, HateCOT facilitates effective K-shot fine-tuning of LLMs with limited data and improves the quality of their explanations, as confirmed by our human evaluation. Our repository is available at https://github.com/hnghiem-usc/hatecot .
HateCOT: An Explanation-Enhanced Dataset for Generalizable Offensive Speech Detection via Large Language Models
Huy Nghiem University of Maryland nghiemh@umd.edu                        Hal Daumé III University of Maryland & Microsoft Research hal3@umd.edu
@@ 1 Introduction
As social media has become indispensable to modern discourse, this channel of communication has amplified the propagation of offensive content. Speech that promotes hateful sentiments thrives on such platforms, leading to real and significant harm on their audience Giachanou and Rosso (2020); Saha et al. (2019). However, “offensive content” is still a contested construct, as what is and is not allowed varies by platform. In research, different approaches analyze semantically similar but still distinct concepts: Cyber-bullying, Toxicity, Sexist, Racist, Hate etc. Poletto et al. (2021); Fortuna et al. (2021); Nghiem et al. (2024); Nguyen et al. (2023), further highlighting this contestedness.
Compounding the challenge, reliable detection of offensive content typically requires significant amounts of data. Sophisticated models tend to be data-hungry, and the process of curating a dataset tailored to a specific use case can be costly, time-consuming, and emotionally challenging for annotators Founta et al. (2018); Toraman et al. (2022). The typical pipeline consists of collecting samples based on topic-relevant key words, then recruiting either crowdworkers or experts to annotate data before developing classification models Paullada et al. (2021). Each step incurs investment and may inject subtle idiosyncrasies proportionate to the size of the downstream dataset, further limiting transferable usefulness to related tasks Fortuna et al. (2021). The size of a dataset also does not necessarily guarantee cross-domain transferrability Poletto et al. (2021); Fortuna et al. (2021).
In practical settings, users may desire transparency from social media platforms. Therefore, the ability to provide human-understandable justification based on platform-specific policy becomes an attractive feature for content moderation. Nevertheless, current techniques often still fail to offer intuitive explanatory signals Yadav et al. (2023); Babaeianjelodar et al. (2022); Ibrahim et al. (2022).
@@
    We release HateCOT (Hate-related Chains-of-Thought), a dataset of over $52,000$ samples consisting of input text, a hate speech label, and an explanation for that label. This corpus is constructed by merging eight datasets with explanations created by using GPT-3.5-Turbo to augment human annotations.
@@
    We demonstrate the benefits of using HateCOT as a pretraining corpus before finetuning on a target domain. Empirical results across 3 datasets show that open-source Large Language Models (LLMs) can effectively leverage definitions to adapt to new tasks using zero-shot and few-shot settings via finetuning and in-context learning .
@@
    We assess the quality of explanations generated by our finetuned models with respect to the criteria described in their corresponding definitions. These insights showcase LLM-generated explanations as a means to enhance transparency in content moderation.
@@ 2.1 Offensive Speech Detection
Cross-domain generalization, however, still remains a relevant challenge in the area. Fortuna et al. (2021) found empirically that cross-dataset transference is highly dependent on semantic similarity between their label spaces. Recent works have pretrained Transformer-based models, such as HateBERT and fBERT, on specialized corpora to enhance generalization to various levels of success Caselli et al. (2021); Sarkar et al. (2021).
@@ 2.2 LLMs in Offensive Speech Classification
Zampieri et al. (2023) assessed a range of open-sourced LLMs on zero-shot prompting on the OffensEval task and found their performance trailing by a wide margin behind trained existing finetuned BERT-based systems. Chiu et al. (2021) and Han and Tang (2022) used the proprietary GPT-3 on a set of different datasets and noted that informative contexts and examples could boost the model’s performance to competitive levels on a different set of data. Similarly, Roy et al. (2023) found that adding explanation to pipeline could result in 10 to 20% boost in performance of LLMs over baselines. (Yang et al., 2023)’s study found that training LLMs with step-by-step reasoning grounded by annotations could improve predictive power.
Pretrained language models have exhibited remarkable ability in text generation Celikyilmaz et al. (2020). Recent large-size LLMs such as GPT-3 and later models are capable of astoundingly fluent, convincing and knowledge-infused outputs Zhang et al. (2023). LLMs with hundreds of billions of parameters even exhibit reasoning capabilities Wei et al. (2022, 2021), leading to a flurry of research on prompting techniques to harness their prowess, such as Chain-of-thought (COT), Tree-of-thought etc. (Yao et al., 2023; Diao et al., 2023). An interesting line of research leverages LLMs to efficiently generate high volumes of synthetic data for tasks with training resource is scarce Puri et al. (2020); Bao et al. (2023); Whitehouse et al. (2023). We build upon these works to construct a dataset that can induce smaller LLMs to efficiently adapt to new categories of offensive content by leveraging their provided definitions.
@@ 3.1 Data Selection
    Target / Rationale: the dataset should provide the a) targets and/or b) rationales, which are fragments of free texts by human annotators to convey some understanding of the corresponding post. Figure 4 shows several examples of rationales, demonstrating their unsuitability to be used as explanations in their native format.
These criteria significantly reduce the number of eligible candidates since many do not provide the required annotation on target and definition. Table 1(a) lists the 8 selected datasets.
@@
Using similar criteria, we select 3 additional datasets with different label spaces and definitions for downstream testing (details in Table 1(b)).
HateCheck was created by Röttger et al. (2021) with the explicit goal of evaluating hate speech detection models. 10 trained annotators labelled the dataset using a binary schema: Hateful and Non-hateful, with the reported inter-annotator agreement coefficient to be 0.93.
HateXplain was primarily collected from Twitter and the Gab platform Mathew et al. (2021). In addition to the labels Hate, Offensive, Normal, annotators also provide justification for their selection by highlighting the span of tokens, called rationales, that contribute to their decision.
Latent_Hate was created on the premise that offensive speech classifiers tend to bias towards covert negative sentiment ElSherief et al. (2021). After discarding the augmented portion from the Social_Bias dataset to avoid contamination, 22,584 samples collected from Twitter remained. This dataset contains 3 coarse-grained labels Not Hate, Explicit Hate, Implicit Hate, while a subset contains 6 fine-grained categories, which we refer to as Implicit_Hate in subsequent test regimens.
@@
Inspired by Yang et al. (2023)’s work that shows that GPT-3.5 could augment human-written rationales to create coherent texts that are still faithful to the original content, we use the prompt template in Figure 1 to generate the explanation, which is guided by the available annotations on label, target, and rationale from the chosen corpora. For datasets that contain multiple annotations per sample, we select the ultimate label via majority voting and concatenate annotations on the targets and/or rationales into a single string delimited by "|".
@@
We use GPT-3.5-Turbo, accessed via the OpenAI’s API, to generate explanations due to this model’s affordability and its capability to follow instructions and generate coherent outputs (Ye et al., 2023a; Koubaa, 2023). For each of the 8 training datasets, we first randomly select and qualitatively analyze 20 samples to ensure the generated explanations are a) stylistically coherent, b) consistent with the provided labels, and c) congruent with the criteria denoted by the definitions. If deemed unsatisfactory, we iteratively adjust the input prompt until the quality threshold is achieved. Appendix A.4 describes this quality assurance process and the final prompts for different scenarios.
@@ 3.2 Optimization of Synthesized Corpus
Extending previous works Magister et al. (2022); Ho et al. (2023), we are interested in optimizing 2 parameters central to the construction of our corpus: the distribution of neutral vs. non-neutral classes in the data and the number of explanations per sample. The former has been noted to influence predictive powers Rathpisey and Adji (2019); Casula and Tonelli (2020). The latter, also referred to as degree of reasoning diversity, could improve knowledge distillation Ho et al. (2023). We use the open-source model Llama 2 Chat-HF of 7 billion parameters (hereby referred to as Llama 7B) (Touvron et al., 2023) to perform tuning experiments in this stage due to its manageable size and strong classification performance. These empirical findings then guide the final augmentation process.
3.2.1 Optimization Procedure
Below are the experiments we perform on a sample of the collected data to optimize these parameters.
@@ 3.2.1 Optimization Procedure
Figure 5 illustrates the Alpaca-styled template to format each post with its corresponding label, generated explanation, definitions along with the instruction into blocks of an input prompt (Taori et al., 2023). Using the described corpus, we supervised finetune Llama 7B via LoRA techniques (technical details specified in Appendix A.2) (Hu et al., 2021). Then, we perform zero-shot classification using the same template to prompt the finetuned model to generate the explanation and label for posts drawn from the test datasets HateCheck, HateXplain, Latent_Hate. We omit Implicit_Hate at this stage due to this set’s markedly different 6-label space. For this part, posts are drawn using K-shot sampling (equal number of samples for each class) on the Validation portion of the test data, based on the values of K val shown in Table 2(b).
@@
For the first experiment, the training data is split into subsets whose distribution between the neutral (NE) class and non-neutral class(es) (NN) described by the following formula: $NN=R*NE$, where $R\in\{1,2,3,4\}$ is the ratio coefficient. We set the number of explanations per sample to 2 , the smallest value that still enables the benefit of reasoning diversity. For the second experiment, we construct the subsets by varying the degree of reasoning diversity $D\in\{1,2,3,4\}$ of each post.
@@
We extract the generated explanation and predicted labels after their respective tags. If the models generate multiple items from the dataset’s label space, we select the first admissible label. If no acceptable output is obtained, we randomly select an item in the label space.
@@ 3.2.2 Insights and Augmentation
We report Llama 7B’s macro F1-scores on the validation set of each configuration in Table 2. A balanced distribution between neutral and non-neutral classes in the training corpus is beneficial, as reflected by the substantially high mean F1-score of 0.643 when R=1. On the other hand, having 3 explanations per sample (D=3) achieves the best overall performance across 3 test sets, consistent with Ho et al. (2023)’s findings on the benefit of multiple reasoning paths. However, performance markedly degrades when D=4. Our manual analysis reveals that the quality of generated outputs deteriorates as the degree of diversity increases, consequently affecting the performance of models trained on this data.
Guided by these empirical findings, we augment the training corpus by selecting approximately 1,800 extra samples from each of the 8 datasets while preserving the 1:1 balanced ratio of neutral to non-neutral classes. Then, we collect 3 explanations per sample using the described mechanism, resulting in a final corpus of 52,137 samples (Table 1(a)), hereby referred to as HateCOT.
@@ 4 Experiments on Test Sets
We perform experiments to answer 3 questions. First, does HateCOT improve zero-shot classification of open-sourced LLMs on unseen datasets? Second, how much data is necessary to enable competitive performance via in-domain finetuning after pretraining on HateCOT? Finally, is in-context learning a viable alternative to finetuning?
@@
    COT-T5-XL A variant of the Flan-T5-XL (3 billion parameters), this model is further finetuned on the CoT dataset, a collection of 1.8 million samples augmented with chain-of-thought-style explanations (Kim et al., 2023).
@@ 4.1 Zero-shot Classification
We prompt the models to perform classification with no in-context examples via 2 modes: No Explanation, where the model directly predicts the label for the input, and With Explanation, where a justification is required before the predicted label. We finetune the base models using only HateCOT and evaluate their performance on the 4 test sets as in Section 3.2.2 (more details in Appendix A.2).
From results presented in Figure 2(a), the smaller models Flan-T5-L and OPT-IML are unable to generate explanations when prompted. In contrast, their scaled-up counterparts could follow instructions at all settings. Asking base (off-the-shelve) models to generate an explanation before the label results in observable boost to Llama models on HateCheck and HateXPlain, but actually hampers performance on Latent_Hate and its derivative Implicit_Hate, which are notably challenging due to its covert nature ElSherief et al. (2021).
@@
Pretraining on HateCOT unanimously enables all models to generate explanations. While smaller models receive no observable boost, larger models (COT-T5-XL, Llama 7B, Llama 13B) are considerably enhanced compared to their base counterparts. With the exception of HateXplain, the HateCOT-pretrained version of COT-T5-XL attains an increment in F1 scores of 7.6% on HateXplain and 9.5% on Latent_Hate over the base counterpart without explanations. Similarly, Llama 7B observes 23.9%, 25.6%, 10% increment on HateCheck, HateXplain and Latent_Hate, respectively. These statistics are 27.9%, 118.5%, and 10.2% for Llama 13B. Notably, all models yield non-competitive performance on Implicit_Hate.
The reduced performance of HateCOT-pretrained models compared to their base counterparts without explanations (e.g.: Flan-T5-L and OPT-IML on almost all test sets) is in line with literature as COT-style prompting tends to favor larger models Wei et al. (2022); Wang et al. (2024); Suzgun et al. (2023). Even the reduced F1 scores of COT-T5-XL on HateXPlain and Implicit_Hate is consistent with this model’s suboptimal performance relative to its larger variants, as showcased in Kim et al. (2023). These results serve as an empirical reference for researchers to select the appropriate model size for their respective task.
(a) Macro F1 scores of LLMs in zero-shot setting using 3 configurations. Base refers to out-of-the-box models, +HateCOT denotes their pretrained counterpart on our dataset. FT5L: Flan-T5-L, CT5XL: COT-T5-XL. Results for Base Flan-T5-L and OPTIML models for With Explanation settings omitted to reflect their inability to generate explanation according to the prompt.
(b) Macro F1 scores for models in zero-shot setting with explanation after K-shot in-domain finetuning at various values of K. Dashed line represents finetuned base models, solid line represents models pre-trained on HateCOT, then in-domain finetuned. For each dataset, the horizontal dashed line represents the base version of the best performing model at K=256 which is finetuned using the entire training data without any rationale for comparison, denoted as Best model (Full data, No. exp).
(c) Min, max and mean of macro F1 scores for Llama 7B and Llama 13B over 5 seed using in-context learning. Pretrained denotes models finetuned on HateCOT only. Finetuned denotes ICL performed on models that have been both HateCOT-pretrained then K=256 shot in-domain finetuned.